# Refining Concentration for Gaussian Quadratic Chaos

### v1

@@ VII Proof of Proposition 5
We study the cases $x>0$ and $x<0$, separately.
@@
    The only positive first-order critical number for $f$ is $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, one easily checks that $f^{\prime}(x)<0$ over an open interval $I$ around 00. Since $f(0)=0$, then the Mean Value Theorem (MVT) implies that $f(x)<0$ over $I\cap(0,\infty)$ and hence, (6) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $10<x_{0}<$ 1, $f^{\prime}(x)>0$ over $(0,x_{0})$ and $f^{\prime}(x)<0$ over $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=-\infty$. This tells us that $f$ rises about 00 on right of $x=0$, reaches an absolute maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to $\infty$- ∞ on left of $x=1$. As such, $f$ possesses a unique zero (x-crossing) somewhere over $(x_{0},1)$. We let $b$ be this zero. Then (6) clearly holds for $0\leq x\leq b$. Finally, we show that $b$ is $\theta_{m}^{-1}(a)$, or equivalently, $\theta_{m}(b)=a$. We have $f(b)=0$, i.e.,
@@
    Recalling the Maclaurin series for $-\ln(1-x)$, we have $-\ln(1-b)=\sum_{j=1}^{\infty}\frac{b^{j}}{j}$. Then (97) can be simplified as
    $\sum_{j=m+1}^{\infty}\frac{b^{j}}{j}=ab^{m+1}$ (98)
@@
    $a=\sum_{j=m+1}^{\infty}\frac{b^{j-m-1}}{j}.$ (99)
    Changing the index $j$ to $j+m+1$, we get $a=\sum_{j=0}^{\infty}\frac{b^{j}}{j+m+1}$, i.e., $a=\theta_{m}(b)$ as promised.
  • Now, let $x<0$. If $m$ is odd, $f(x)$ is given by $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ as in the previous case and one easily checks that for $a>\frac{1}{n+1}$, $f^{\prime}(x)$ given in (96) is negative for every $x<0$. Then MVT implies that $f(x)>0$ for every $x<0$ as desired. If $m$ is even, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}-ax^{m+1}+\ln(1-x)$. A similar computation as in (96) gives that
@@
    Since $m$ is even, once again we have $f^{\prime}(x)<0$ for every $x>0$ and MVT comes in one last time to confirm that $f(x)>0$ for $x<0$.
    VIII Appendix B; Proof of Lemma 2
@@
    $f^{\prime}(x)=\frac{x^{m+1}((m+2)a-m-1-a((m+1)a-m)x)}{(m+1)(1-x)(1-ax)^{2}}.$ (102)
    The only positive first-order critical number for $f$ over the interval $(0,1)$ is given by $x_{0}=\frac{m+1-(m+2)a}{a(m-(m+1)a)}$. Moreover, $f^{\prime}(x)>0$ for $0<x<x_{0}$ and $f^{\prime}(x)<0$ for $x_{0}<x<1$. Since $f(0)=0$, then MVT implies that $f(x)>0$ for all $x$ in the interval $(0,x_{0}]$ and hence, (10) holds with $b=x_{0}$.
Appendix C; The choice of $a=a_{0}$ in (16) results in the largest $\kappa$
In order to compare $f(s_{opt})$ in (53) with $\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}$, we address the cases $c\leq 1$ and $c>1$ separately.
@@
    $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq c\\ \frac{b}{2\rho}-\frac{b^{2}}{4(1-ab)\rho^{2}}&c<\rho\leq 1\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>1\end{array}\right..$ (106)
@@
    $-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq 1\\ \frac{1}{2a}-\frac{1}{2a^{2}\rho}(\sqrt{1+2a\rho}-1)&1<\rho\leq c\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>c\end{array}\right..$ (110)
    It is routine to check that this is a continuous function of $a$ and it achieves its absolute minimum value of $\frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)$ at $\rho=1$.
@@
$\min_{\rho>0}-\frac{f(s_{opt})}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}&\frac{2}{3}<a\leq a_{0}\\ \frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)&a_{0}<a<1\end{array}\right..$ (113)
We need to choose $\frac{2}{3}<a<1$ such that the right side of (113) is as large as possible. Fig. LABEL:pic_1111 shows the graph of this function over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
Figure 3: The graph of the function of $a$ on the right side of (113) over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$
@@
Define the function
@@
Tedious algebra shows that
$-ux+vx^{2}+wx^{3}=\frac{g(u,v,w)}{27w^{\frac{1}{2}}}.$ (115)
@@
Then $\frac{\partial g}{\partial w}\leq 0$ follows by noting that $2v^{2}+3uw\geq 2v\sqrt{v^{2}+3uw}$ for all $u,v,w$.

### v2

@@ X Concluding Remarks
We study the cases $x>0$ and $x<0$ separately.
  1. 1.
@@
    The only first-order critical numbers for $f$ are 00 and $x_{0}=1-\frac{1}{(m+1)a}$. If $a\leq\frac{1}{m+1}$, then one easily checks that $f^{\prime}(x)<0$ over the open interval $(0,1)$. Since $f(0)=0$, the Mean Value Theorem (MVT) implies that $f(x)<0$ over $(0,1)$ and hence, (7) can not hold for any $b>0$. Conversely, let $a>\frac{1}{m+1}$. Then $10<x_{0}<$ 1, $f^{\prime}(x)>0$ over the interval $(0,x_{0})$ and $f^{\prime}(x)<0$ over the interval $(x_{0},1)$. Note that $\lim_{x\to 1^{-}}f(x)=-\infty$. This tells us that $f$ rises above 00 on right of $x=0$, reaches a maximum value at the critical number $x_{0}$ and then goes down from there on and escapes to $\infty$- ∞ on left of $x=1$. As such, $f$ has a unique zero (x-crossing) somewhere over the interval $(x_{0},1)$. We let $b$ be this zero. Then (7) clearly holds for $0\leq x\leq b$. Finally, we show that $a=\theta_{m}(b)$. We have $f(b)=0$, i.e.,
@@
    Recalling the Maclaurin series for $-\ln(1-x)$, we have $-\ln(1-b)=\sum_{i=1}^{\infty}\frac{b^{i}}{i}$. Then (209) is simplified as
    $\sum_{i=m+1}^{\infty}\frac{b^{i}}{i}=ab^{m+1}$ (210)
@@
    $a=\sum_{i=m+1}^{\infty}\frac{b^{i-m-1}}{i}.$ (211)
    Changing the index $i$ to $i+m+1$, we get $a=\sum_{i=0}^{\infty}\frac{b^{i}}{i+m+1}$, i.e., $a=\theta_{m}(b)$ as promised.
  2. 2.
    Let $x<0$. If $m$ is odd, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}+ax^{m+1}+\ln(1-x)$ as in the previous case and one easily checks that for $a>\frac{1}{m+1}$, $f^{\prime}(x)$ given in (208) is negative for every $x<0$. Then MVT rules that $f(x)>0$ for every $x<0$ as desired. If $m$ is even, then $f(x)=x+\frac{x^{2}}{2}+\frac{x^{3}}{3}+\cdots+\frac{x^{m}}{m}-ax^{m+1}+\ln(1-x)$. A similar computation as in (208) gives that
@@
    Since $m$ is even, once again we have $f^{\prime}(x)<0$ for every $x<0$ and MVT enters once again to confirm that $f(x)>0$ for $x<0$.
Appendix B; Proof of Lemma 2
@@
$f^{\prime}(x)=\frac{x^{m+1}\big{(}(m+2)a-m-1-a((m+1)a-m)x\big{)}}{(m+1)(1-x)(1-ax)^{2}}.$ (214)
Thanks to the condition $\frac{m+1}{m+2}<a<1$, the only positive first-order critical number for $f$ over the interval $(0,1)$ is given by $x_{0}=\frac{m+1-(m+2)a}{a(m-(m+1)a)}$. Moreover, $f^{\prime}(x)>0$ for $0<x<x_{0}$ and $f^{\prime}(x)<0$ for $x_{0}<x<1$. Since $f(0)=0$, then MVT implies that $f(x)>0$ for every $x$ in the interval [^21] In fact, $f$ has a unique x-crossing at some number $x_{1}$ inside the interval $(x_{0},1)$ and $f(x)\geq 0$ for every $0\leq x\leq x_{1}$. Hence, (120) also holds with $b=x_{1}$. However, unlike $x_{0}$, the number $x_{1}$ does not admit a closed-form expression. $(0,x_{0}]$ and hence, (120) holds with $b=x_{0}$.
Appendix C; The choice of $a=a_{0}$ in (129) yields the largest $\kappa$
In order to compare $\Lambda(t,a)$ in (17) with $\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}$, we address the cases $c\leq 1$ and $c>1$ separately.
@@
    $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq c\\ \frac{b}{2\rho}-\frac{b^{2}}{4(1-ab)\rho^{2}}&c<\rho\leq 1\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>1\end{array}\right..$ (218)
@@
    $-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{1}{2a\rho}-\frac{1}{2a^{2}\rho^{2}}\big{(}\sqrt{1+2a\rho}-1\big{)}&0<\rho\leq 1\\ \frac{1}{2a}-\frac{1}{2a^{2}\rho}(\sqrt{1+2a\rho}-1)&1<\rho\leq c\\ \frac{b}{2}-\frac{b^{2}}{4(1-ab)\rho}&\rho>c\end{array}\right..$ (222)
    Once again, this is a continuous function of $\rho$ and it achieves its absolute minimum value of $\frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)$ at $\rho=1$.
@@
$\min_{\rho>0}-\frac{\Lambda(t,a)}{\min\{\frac{t^{2}}{\beta^{2}},\frac{t}{\alpha}\}}=\left\{\begin{array}[]{cc}\frac{6a^{3}-a^{2}-8a+4}{4a^{2}(2a^{2}-3a+1)}&\frac{2}{3}<a\leq a_{0}\\ \frac{1}{2a}-\frac{1}{2a^{2}}(\sqrt{1+2a}-1)&a_{0}<a<1\end{array}\right..$ (225)
We need to choose $\frac{2}{3}<a<1$ such that the right hand side in (225) is as large as possible. Fig. 13 shows the graph of this function of $a$ over the interval $(\frac{2}{3},1)$. We see that it achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
Figure 13: The graph of the function of $a$ that sits on the right side of (225) over the interval $(\frac{2}{3},1)$. It achieves its absolute maximum value of $\frac{9-\sqrt{17}}{32}$ at $a=a_{0}=\frac{7-\sqrt{17}}{4}$.
@@
Tedious algebra shows that
$-ux+vx^{2}+wx^{3}=\frac{g(u,v,w)}{27w^{\frac{1}{2}}},$ (226)
where $g(u,v,w)$ is defined by
@@
Then $\frac{\partial g}{\partial w}\leq 0$ follows by noting that $2v^{2}+3uw\geq 2v\sqrt{v^{2}+3uw}$ for all $u,v,w>0$.
Appendix E; Proof of (32)
For an even integer $n\geq 8$, let us denote $m=\frac{n}{2}\geq 4$. During the proof, we adopt the notation $\forall_{\infty}m$ to say that a statement indexed by the integer $m$ holds for all but finitely many values for $m$, i.e., it eventually holds for all sufficiently large $m$. We begin by showing that
$\forall_{\infty}m\geq 4,\,\,\,r^{\prime}_{2m}>2m.$ (229)
This establishes the second claim in (32). Recall from the proof of Proposition 7 that $r_{n},r^{\prime}_{n}$ are the only two positive zeros for the polynomial $P_{n}^{(1)}(r)-P_{n}^{(2)}(r)$ and that this polynomial is negative if and only if $r_{n}<r<r^{\prime}_{n}$. In order to verify (229), it is enough to show that
$\forall_{\infty}m\geq 4,\,\,\,P_{2m}^{(1)}(2m)<P_{2m}^{(2)}(2m).$ (230)
We have $P_{2m}^{(2)}(2m)=\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{1}{2^{i}i!}(1+2m)^{i}$. We show that the last term in this sum is eventually larger than $P_{2m}^{(1)}(2m)=2^{m}$, i.e.,
[…]
This is easily done by invoking the Stirling’s approximation formula. Let us recall the inequality [^22] See Section 2.9 in [8].
$k!<\sqrt{2\pi}\,k^{k+\frac{1}{2}}e^{-k+\frac{1}{12k}},\,\,\,k\geq 1,$ (232)
due to Feller [8, 9]. Plugging $k=m-1\geq 3$ in (232) and observing that $\sqrt{2\pi}e^{\frac{1}{12(m-1)}}\leq\sqrt{2\pi}e^{\frac{1}{36}}\approx 2.577<e$, we get $(m-1)!<(m-1)^{m-\frac{1}{2}}e^{2-m}$ and hence,
$\frac{e^{m-2}}{(m-1)^{m-\frac{1}{2}}}<\frac{1}{(m-1)!},\,\,\,m\geq 4.$ (233)
By (233), the proof of (231) is complete if we can verify the stronger statement
[…]
Note that $\frac{(m+\frac{1}{2})^{m-1}}{(m-1)^{m-\frac{1}{2}}}>\frac{(m+\frac{1}{2})^{m-1}}{(m+\frac{1}{2})^{m-\frac{1}{2}}}=\frac{1}{(m+\frac{1}{2})^{\frac{1}{2}}}$. Thus, (234) follows if
[…]
But, this is certainly true as the right hand side scales like $e^{m-O(\frac{1}{2}\ln m)}$. The proof of (229) is now complete.
To verify the first claim in (32), we show that
$\forall\epsilon>0,\,\,\,\forall_{\infty}m\geq 4,\,\,\,r_{2m}<\epsilon.$ (236)
Following the discussion before (230), it suffices to prove that
$\forall_{\infty}m\geq 4,\,\,\,P_{2m}^{(1)}(\epsilon)<P_{2m}^{(2)}(\epsilon).$ (237)
The proof follows similar lines of reasoning to show $P_{n}^{(1)}(1)<P_{n}^{(2)}(1)$ in Lemma 4. Fix $\epsilon>0$. Let $\boldsymbol{x}\sim\mathrm{Poisson}(\frac{1+\epsilon}{2})$ be a Poisson random variable with parameter $\frac{1+\epsilon}{2}$. Then
$P_{2m}^{(2)}(\epsilon)$ $=$ $\frac{1}{\sqrt{e}}\sum_{i=0}^{m-1}\frac{(1+\epsilon)^{i}}{2^{i}i!}$ (238)
$=$ $e^{\frac{\epsilon}{2}}\sum_{i=0}^{m-1}\frac{e^{-\frac{1+\epsilon}{2}}(\frac{1+\epsilon}{2})^{i}}{i!}$
$=$ $e^{\frac{\epsilon}{2}}\Pr(\boldsymbol{x}\leq m-1)$
$=$ $e^{\frac{\epsilon}{2}}(1-\Pr(\boldsymbol{x}\geq m))$
$\stackrel{{\scriptstyle(a)}}{{\geq}}$ $e^{\frac{\epsilon}{2}}\Big{(}1-\frac{\mathbb{E}[\boldsymbol{x}^{2}]}{m^{2}}\Big{)}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $e^{\frac{\epsilon}{2}}\Big{(}1-\frac{\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}}{m^{2}}\Big{)}$
$=$ $e^{\frac{\epsilon}{2}}\Big{(}1-\frac{\epsilon^{2}+4\epsilon+3}{4m^{2}}\Big{)},$
where $(a)$ follows by Markov’s inequality and $(b)$ is due to $\mathbb{E}[\boldsymbol{x}^{2}]=\frac{1+\epsilon}{2}+(\frac{1+\epsilon}{2})^{2}$. Also, $P_{2m}^{(1)}(\epsilon)=(1+\frac{\epsilon}{2m})^{m}$. By (238), if we can show that
$\forall_{\infty}m\geq 4,\,\,\,\Big{(}1+\frac{\epsilon}{2m}\Big{)}^{m}<e^{\frac{\epsilon}{2}}\Big{(}1-\frac{\epsilon^{2}+4\epsilon+3}{4m^{2}}\Big{)},$ (239)
then (237) follows. Define the function
$f_{\epsilon}(x)=\frac{\epsilon}{2}+\ln\Big{(}1-\frac{\epsilon^{2}+4\epsilon+3}{4x^{2}}\Big{)}-x\ln\Big{(}1+\frac{\epsilon}{2x}\Big{)},\,\,\,x>\frac{1}{2}\sqrt{\epsilon^{2}+4\epsilon+3}.$ (240)
The statement in (239) can be written as $\forall_{\infty}m\geq 4,e^{f_{\epsilon}(m)}>1$. Thus, it is enough to prove that $f_{\epsilon}(x)>0$ for sufficiently large $x$. This is accomplished if we can verify
$\lim_{x\to\infty}f_{\epsilon}(x)=0$ (241)
and
$f^{\prime}_{\epsilon}(x)<0,$ (242)
for all sufficiently large $x$. Checking (241) is trivial. To verify (242), we use a similar argument and show that
$\lim_{x\to\infty}f^{\prime}_{\epsilon}(x)=0$ (243)
and
$f^{\prime\prime}_{\epsilon}(x)>0,$ (244)
for all sufficiently large $x$. We have
$f^{\prime}_{\epsilon}(x)=\frac{2(\epsilon^{2}+4\epsilon+3)}{x(4x^{2}-\epsilon^{2}-4\epsilon-3)}+\frac{\epsilon}{2x+\epsilon}-\ln\Big{(}1+\frac{\epsilon}{2x}\Big{)}.$ (245)
Then (243) clearly holds. Also,
$f^{\prime\prime}_{\epsilon}(x)=\frac{p_{\epsilon}(x)}{x^{2}(2x+\epsilon)^{2}(4x^{2}-\epsilon^{2}-4\epsilon-3)^{2}},$ (246)
where $p_{\epsilon}(x)$ is a polynomial of degree $5$ in $x$ given by [^23] The coefficients behind powers of $x$ in $p_{\epsilon}(x)$ are all polynomials in $\epsilon$.
$p_{\epsilon}(x)=16\epsilon^{2}x^{5}+\textrm{terms of lower degrees in $x$}.$ (247)
The leading term in the polynomial $p_{\epsilon}(x)$ is $16\epsilon^{2}x^{5}$. Since the coefficient $16\epsilon^{2}$ behind $x^{5}$ is positive, then (244) must hold for all sufficiently large $x$ depending on $\epsilon$ of course. The proof of (236) is complete.